# https://github.com/NVlabs/OmniDrive

### v1

@@ Abstract
The advances in multimodal large language models (MLLMs) have led to growing interests in LLM-based autonomous driving to leverage their strong reasoning capabilities. However, capitalizing on MLLMs’ strong reasoning capabilities for improved planning behavior is challenging since it requires full 3D situational awareness beyond 2D reasoning. To address this challenge, our work proposes OmniDrive, a holistic framework for strong alignment between agent models and 3D driving tasks. Our framework starts with a novel 3D MLLM architecture that uses sparse queries to lift and compress visual representations into 3D before feeding them into an LLM. This query-based representation allows us to jointly encode dynamic objects and static map elements (e.g., traffic lanes), providing a condensed world model for perception-action alignment in 3D. We further propose a new benchmark with comprehensive visual question-answering (VQA) tasks, including scene description, traffic regulation, 3D grounding, counterfactual reasoning, decision making and planning. Extensive studies show the excellent reasoning and planning capabilities of OmniDrive in complex 3D scenes.
Keywords:
Autonomous driving Large language model Planning
Figure 1: We present OmniDrive, a novel framework for end-to-end autonomous driving with large language model (LLM) agents. Our main contributions involve novel solutions in both model (OmniDrive-Agent) and benchmark (OmniDrive-nuScenes). The former features a novel 3D vision-language model design, whereas the latter is constituted of comprehensive VQA tasks for reasoning and planning.
@@ 1 Introduction
The recent rapid development of multimodal LLMs (MLLMs) [1, 24, 31] and their excellent reasoning capabilities have led to a stream of applications in end-to-end autonomous driving [46, 58, 52, 39, 6]. However, the challenge to extend the capabilities from 2D understanding to the intricacies of 3D space is a crucial hurdle to overcome to fully unlock its potential in real-world applications. Understanding and navigating through 3D space is indispensable for autonomous vehicles (AVs) because they directly impact an AV’s ability to make informed decisions, anticipate future states, and interact safely with their environment. Although previous works [38, 47] have demonstrated successful applications of LLM-agents in autonomous driving, a holistic and principled approach is still needed to fully extend the 2D understanding and reasoning capabilities of MLLMs into complex 3D scenes for understanding the 3D geometry and spatial relations.
Another open issue is the need to address multi-view high resolution video input. On one hand, many current popular 2D MLLM architectures, such as LLaVA-1.5 [31, 30], can only take $336\times 336$ image input due to the limited vision encoder resolution and LLM token sequence length. Increasing the limitations is not trivial, as it requires significantly more compute and memories. On the other hand, dealing with high resolution video input, oftentimes even multi-view, is a fundamental requirement for long-range AV perception and safe decision making. However, unlike many cloud-based services, real-world industrial autonomous driving applications are mostly on-device and compute-bound. As such, there is a need to design an efficient MLLM architectures with compressed 3D visual representations before feeding to the LLM.
Our answer to the above challenges is a novel Q-Former-styled [24] 3D MLLM architecture as shown in  Fig. 1. Unlike LLaVA which adopts a self-attention design, the cross-attention decoder in Q-Former makes it more scalable to higher resolution input by compressing the visual information into sparse queries. Interestingly, we notice that the Q-Former architecture shares considerable similarity with the family of perspective-view models, such as DETR3D [53], PETR(v2) [32, 33], StreamPETR [50] and Far3D [21]. Using sparse 3D queries, these models have demonstrated considerable advantages over dense bird’s-eye view (BEV) representation with leading performance [50, 21], long-range perception [21] and capability to jointly model map elements [55]. The similarity in query-based decoder architecture enables us to align both worlds by appending 3D position encoding to the queries, lifting them to 3D, and attending the multi-view input, as shown in the left portion of Fig. 1. This process allows the MLLM to gain 3D spatial understanding with minimal efforts and changes, while leveraging the pre-trained knowledge from the abundant images in 2D.
@@
Besides model architectures, recent drive LLM-agent works also feature the importance of benchmarks [43, 38, 46, 39, 10, 47]. Many of them are presented as question-answering (QA) datasets to train and benchmark the LLM-agent for either reasoning or planning. Despite the various QA setups, benchmarks that involve planning [46, 10, 47] still resort to an open-loop setting on real-world sessions (e.g., nuScenes) where expert trajectories are used. Recent studies [60, 26], however, reveal limitations of open-loop evaluation with implicit biases towards ego status, overly simple planning scenarios, and easy overfit to the expert trajectory. In light of the issue, the proposed benchmark OmniDrive-nuScenes features a counterfactual reasoning benchmark setting where simulated decision and trajectories are leveraged to reason potential consequences. Our benchmark also include other challenging tasks that require full 3D spatial understanding and long-horizon reasoning, as shown in Fig. 1 (right).
In summary, OmniDrive aims to provide a holistic framework for end-to-end autonomous driving with LLM-agents. We propose a principled model design excellent in 3D reasoning and planning, as well as a more challenging benchmark going beyond single expert trajectories.
2 OmniDrive-Agent
As a recap, we aim for a unified 3D MLLM design to: 1) leverage the 2D MLLM pre-training knowledge, and 2) addressing the high-resolution multi-view input in autonomous driving. We propose a Q-Former-styled architecture by compressing the visual features into a fixed number of queries before feeding to an LLM [24]. Noticing the similarity between Q-Former and query-based 3D perception frameworks, we align our MLLM architecture with StreamPETR [50], where use queries to encode both dynamic objects and static map elements. These queries, together with additional carrier tokens, serve as a condensed world model to align perception with reasoning and planning.
2.1 Preliminaries
The Q-Former based MLLMs are composed of a general visual encoder to extract single-view image features $F_{s}\in\mathbb{R}^{C\times H\times W}$, a projector (Q-Former) that serves as visual-language alignment module, and a large language model for text generation. The architecture of the projector is stacked transformer decoder layers. The projection process from image features to the textual embedding can be represented as:
$\tilde{Q}_{t}=f_{q}(Q_{t},{F}_{s})$ (1)
where $Q_{t}$ is the initialized text embedding. $\tilde{Q}_{t}$ is the refined text embedding, which will be sent to the language model to generate the final text output.
The query-based 3D perception models [55, 21, 28, 29] consist of a shared visual encoder to extract multi-view image features, and a detection head $f_{d}$. It is based on the PETR [32] and utilizes transformer decoder architecture to efficiently convert multi-view image features $F_{m}\in\mathbb{R}^{N\times C\times H\times W}$ into detection queries $\tilde{Q}_{d}$, which can be formulated as:
$\tilde{Q}_{d}=f_{d}(Q_{d},F_{m}+P_{m})$ (2)
where $P_{m}$ is the 3D position encoding that effectively capturing the geometric relationship between the image view and 3D domains. $Q_{d}$ is the initialized detection queries to gather the multi-view image features.
Figure 2: Overall pipeline of OmniDrive-Agent. The left diagram illustrates the overall framework of the model. We employ a 3D perception task to guide Q-Former’s learning. The right diagram depicts the specific structure of Q-Former3D, which is consist of six transformer decoder layers. The attention weights are initialized from 2D pre-pretrain. The input are multi-view image features. Additionally, 3D position encoding is added in the attention operation. Furthermore, we introduce temporal modeling through a memory bank.
It can be observed that the Transformer decoder in Q-Former and the sparse query-based 3D perception models, represented by StreamPETR [50], share highly similar architecture designs. To enhance the localization abilities of the MLLMs, we consider introducing the design of 3D position encoding and the supervision of the query-based perception models.
2.2 Overall Architecture
As show in Fig. 2, Omnidrive first uses a shared visual encoder to extract multi-view image features $F_{m}\in\mathbb{R}^{N\times C\times H\times W}$. The extracted features with the position encoding $P_{m}$ are fed into the Q-Former3D. In Q-Former3D, we initialize the detection queries and carrier queries and perform self-attention to exchange their information, which can be summarized by the following formula:
$\tilde{Q}=\text{Multi-head Attention}(Q,K,V)$
[…]
$(Q,K,V)=(\textbf{[}Q_{c},{Q}_{d}\textbf{]},P_{m}+F_{m},F_{m}),$ (4)
$\tilde{Q}=\text{Multi-head Attention}(Q,K,V)$
After that, the perception queries are used to predict the categories and coordinates of the foreground elements. The carrier queries are sent to a single layer MLP to align with the dimension of LLM tokens (e.g., 4096 dimensions in LLaMA [48]) and further used for text generation following LLaVA [31].
In our model, the carrier queries play the role of the visual-language alignment. Additionally, this design enables carrier queries to leverage the geometric priors provided by the 3D position encoding, while also allowing them to leverage query-based representations acquired through the 3D perception tasks.
2.3 Multi-task and Temporal Modeling
Our approach benefits from multi-task learning and temporal modeling [33, 25]. In multi-task learning, we can integrate task-specific Q-Former3D modules for each perception task, employing a uniform initialization strategy (please refer to Sec. 2.4). In different tasks, carrier queries can gather information of different traffic elements. In our implementation, we cover tasks such as center-line construction and 3D object detection. During the training and inference phases, both heads share the same 3D position encoding. Regarding temporal modeling, we store the perception queries with top-k classification scores into a memory bank and propagate them frame by frame [28, 59]. The propagated queries interact with the perception and carrier queries from the current frame through cross-attention, expanding the capabilities of our model to effectively process video input.
2.4 Training Strategy
The training of OmniDrive-agent comprises two stages: 2D-Pretraining and 3D-Finetuning. In the initial stage, we pretrain the MLLMs on 2D image tasks to initialize the Q-Former and carrier queries. Following this, the model is fine-tuned on 3D-related driving tasks (e.g., motion planning, 3D grounding, etc.). In both stages, we calculate the text generation loss without considering contrasting learning and matching loss for in BLIP-2 [24].
2D-Pretraining. The 2D-Pretraining stage aims to pretrain the carrier queries and the Q-Former, and achieve better alignment between image features and the large language model. When removing the detection queries, the OmniDrive model can be viewed as a standard vision language model capable of generating text conditioned on images. Therefore, we adopt the training recipe and data from LLaVA v1.5 [30] to pretrain OmniDrive on 2D images. The MLLMs is first trained on 558K image-text pairs, during which all parameters except the Q-Former are frozen. Subsequently, we fine-tune the MLLMs using the instruction tuning dataset from LLaVA v1.5. During this fine-tuning step, only the image encoder is frozen, while all other parameters are trainable.
3D-Finetuning. During the 3D finetuning phase, we aim to enhance the model’s 3D localization capability while retaining its 2D semantic understanding as much as possible. We have augmented the original Q-Former with 3D position encoding and temporal modules. In this phase, we fine-tune the visual encoder and the large language model with Lora [16] using a small learning rate, and train Q-Former3D with a relatively larger learning rate.
3 OmniDrive-nuScenes
@@
To benchmark drive LLM-agents, we propose OmniDrive-nuScenes, a novel benchmark built on nuScenes [4] with high quality visual question-answering (QA) pairs covering perception, reasoning and planning in 3D domain.
OmniDrive-nuScenes features a fully-automated procedural QA generation pipeline using GPT4. Similar to LLaVA [31], the proposed pipeline takes the 3D perception ground truths as context information via prompt input. Traffic rules and planning simulations are further leveraged as additional inputs, thereby easing the challenges GPT-4V faces in comprehending 3D environments. Our benchmark asks long-horizon questions in the forms of attention, counterfactual reasoning, and open-loop planning. These questions challenge the true spatial understanding and planning capabilities in 3D space as they require planning simulations in the next few seconds to obtain the correct answers.
Besides using the above pipeline to curate the offline question-answering sessions for OmniDrive-nuScenes, we further propose a pipeline to online generate various types of grounding questions. This part can also be viewed as certain form of implicit data augmentation to enhance the 3D spatial understanding and reasoning capabilities of the models.
3.1 Offline Question-Answering.
@@
The images depict a daytime setting in a controlled-access area, likely a parking lot or a service entrance of a commercial or industrial facility. On the left, there’s a grassy area with trees and a building with blue accents. Moving towards the center, we see a security checkpoint with a raised barrier arm, indicating an entrance or exit point, and a security guard booth…<omitted>
Prompt type 2: Lane-object association
|— your current straight lane [(-2.6, +0.5), (+1.2, +0.7), (+5.0, +0.9), (+8.8, +1.0)]
| |— your own car
| |— movable_object.trafficcone in the front location: (+8.2, +2.4)…<omitted>
Prompt type 3: Simulated decision and trajectory
@@
Prompt type 4: Expert decision and trajectory
@@
Objects near your path: vehicle.moving at (-6.0, +0.3), human.pedestrian.moving at (+4.4, -7.2).
@@
Answer: The most suitable trajectory to follow would be similar to the one provided as safe, which is [PT, (+0.76, +0.02), (+1.45, +0.03),…,(+3.44, +0.12)]. This trajectory ensures that you remain in your current straight lane [(-2.6, +0.5), (+1.2, +0.7), …], avoid any obstacles…<omitted>
@@
Table 1: An example to illustrate the instruction-following data generation for offline QAs. The top block shows contexts such as captions and boxes used to prompt GPT4, and the bottom block shows the four types of responses. Note that the visual image is not used to prompt GPT4 except for caption generation.
Tab. 1 shows an example of the proposed offline data generation pipeline, where in-context information is used to generate the QA pairs on nuScenes. We begin with related details on how different types of the prompt information is obtained:
Caption. When both the image and lengthy scene information are fed into GPT-4V simultaneously, GPT-4V tends to overlook details in the image. So we first prompt GPT-4V to produce the scene description based on the multi-view input only. As shown in the top block of Tab. 1, we stitch the three frontal views and three rear views into two separate images, and feed them into GPT-4V, as shown in the top of the Tab. 1. We also prompt GPT-4V to include the following details: 1) mentions weather, time of day, scene type and other image contents; 2) understands the general direction of each view (e.g., the first frontal view being front-left); 3) avoids mentioning the contents from each view independently and replaces with positions relative to the ego vehicle.
Lane-object association. For GPT-4V, understanding the relative spatial relationships of traffic elements (such as objects, lane lines, etc.) in the 3D world is a highly challenging task. Directly inputting the 3D object coordinates and the curve representation of lane lines to GPT-4V does not enable effective reasoning. Therefore, we represent the relationships between objects and lane lines in the form of a file tree, and convert the information of objects into a natural language description based on their 3D bounding boxes.
Simulated trajectories. Trajectories are sampled for counterfactual reasoning in two ways: 1) We select the initial lane based on three driving intentions: lane keeping, left lane change, and right lane change. Then we use the Depth-First Search (DFS) algorithm to link the lane centerline and get all possible vehicle trajectory paths. Then selecting different completion rates and speed targets for various lanes (acceleration, deceleration, and speed maintenance) to create simulated trajectories. 2) Generating driving trajectories based solely on lane centerlines makes it difficult to simulate scenarios that are ‘out of the drivable’. Therefore, we also performed clustering on the entire nuScenes dataset’s ego trajectories, selecting representative driving paths each time.
Expert trajectory. This is the log replay trajectory from nuScenes. The expert trajectories are classified into different types for high-level decision making. We also identify an object as “close”, if its minimum distance to the trajectory is smaller than 10 meters in the next 3 seconds. The close objects are then listed below the expert trajectory.
In the bottom block of the Tab. 1, we describe the different types of QA responses obtained by using the above context information:
@@
Attention. Given the simulated and expert trajectories, run simulation to identify close objects. At the same time, we also allowed GPT4 to use its own common sense to identify threatening traffic elements.
@@
Decision making and planning. We present the high-level decision making as well as the expert trajectory and use GPT-4V to reason why this trajectory is safe, given the previous prompt and response information as context.
@@
3.2 Online Question-Answering.
Figure 3: Online QA generation.
To fully leverage the 3D perception labels in the autonomous driving dataset, we generate numerous grounding-like tasks during the training process in an online manner. Specifically, the following tasks (in Fig. 3) are designed:

### v2

# OmniDrive: A Holistic Vision-Language Dataset for Autonomous Driving with Counterfactual Reasoning

@@ Abstract
The advances in vision-language models (VLMs) have led to a growing interest in autonomous driving to leverage their strong reasoning capabilities. However, extending these capabilities from 2D to full 3D understanding is crucial for real-world applications. To address this challenge, we propose OmniDrive, a holistic vision-language dataset that aligns agent models with 3D driving tasks through counterfactual reasoning. This approach enhances decision-making by evaluating potential scenarios and their outcomes, similar to human drivers considering alternative actions. Our counterfactual-based synthetic data annotation process generates large-scale, high-quality datasets, providing denser supervision signals that bridge planning trajectories and language-based reasoning. Futher, we explore two advanced OmniDrive-Agent frameworks, namely Omni-L and Omni-Q, to assess the importance of vision-language alignment versus 3D perception, revealing critical insights into designing effective LLM-agents. Significant improvements on the DriveLM Q&A benchmark and nuScenes open-loop planning demonstrate the effectiveness of our dataset and methods.
Figure 1: OmniDrive is a holistic vision-language dataset for autonomous driving, utilizing counterfactual reasoning to generate high-quality QA data from simulated and actual trajectories. We explore two baseline models: Omni-Q, which designs vision-language models (VLMs) from a 3D perception standpoint, and Omni-L, which builds from VLMs to enhance 3D integration.
@@ 1 Introduction
The recent rapid development of 2D Vision-Language Models (VLMs) [1, 20, 25] and their strong reasoning capabilities have led to a stream of applications in end-to-end autonomous driving [41, 53, 48, 34, 4]. However, extending capabilities from 2D to 3D understanding is crucial for unlocking potential in real-world applications. Although previous works [33, 42] have shown successful applications of LLM-agents in autonomous driving (AD), a holistic and principled framework from dataset to LLM-agent is needed to fully extend VLMs’ 2D understanding and reasoning capabilities to 3D geometric and spatial understanding.
{NoHyper}**footnotetext: Work done during an internship at NVIDIA.footnotetext: Corresponding author: zhidingy@nvidia.com.
@@
Recent drive LLM-agent works feature the importance of datasets [38, 33, 41, 34, 8, 42]. Many are presented as question-answering (Q&A) datasets to train and benchmark the LLM-agent for either reasoning or planning. Noteably, benchmarks that involve planning [41, 8, 42] still resort to using expert trajectories for an open-loop setting on real-world sessions (e.g. nuScenes). However, recent studies [54, 22] reveal several limitations of open-loop evaluation: implicit biases towards ego status, overly simple planning scenarios, and easy overfit to expert trajectories.
OmniDrive: VLM Dataset. Expert driving actions provide only sparse supervision [5, 23], primarily reflecting safe trajectories without delving into the complex decision-making and underlying reasoning processes. Relying solely on this sparse supervision makes it challenging to effectively optimize end-to-end driving models. Counterfactual reasoning involves evaluating potential scenarios and their outcomes, similar to how human drivers consider various possibilities to make safer decisions. Therefore, we combine counterfactual reasoning with the chain-of-thought capabilities of VLMs, as shown in Fig. 1. This approach creates a more effective connection between planning trajectories and language-based reasoning.
Additionally, we found that using simulated trajectories for counterfactual reasoning efficiently identifies key traffic elements in a scene. This process creates a structured and simplified 3D scene representation, making it easier for GPT to understand 3D scenes and generate more effective 3D driving Q&A data. To ensure quality, we utilize a rule-based checklist to assess the consequences of potential trajectories. Based on these results, we design prompts for GPT-4 to generate coherent Q&A, which helps identify which objects require attention and evaluate outcomes based on trajectories. A human-in-the-loop approach is employed in designing the checklist and prompts, ensuring comprehensive coverage of all scenarios. This methodology ensures that data generation is both reliable and interpretable.
Omni-L/Q: LLM-agents. Designing effective Driving Vision-Language Models [41, 8, 48] (VLMs) presents a complex and underexplored challenge. A fundamental question is whether to build upon existing 2D VLMs[25, 26] and align them with 3D space, or to integrate current 3D perception stacks [21, 49, 27, 35] into a vision-language framework. To address this, we explore two promising Large Language Model (LLM) frameworks Omni-L and Omni-Q. The Omni-L utilizes state-of-the-art (SoTA) MLP-projection approach (LLaVA [25, 24]), which enhances the performance of existing VLMs. The Omni-Q is based on the BEV architecture employed by StreamPETR [27, 28, 46], incorporating Q-Former’s [20] design to investigate the synergy between LLMs and traditional autonomous driving perception tasks. By addressing crucial considerations in the design of autonomous driving LLM-Agents, we conduct a comprehensive comparison of these paradigms in tasks such as counterfactual reasoning and open-loop planning. Our findings indicate that migrating 2D VLMs to 3D is a more straightforward approach compared to integrating traditional 3D perception stacks into VLMs.
Contributions. We propose OmniDrive, a holistic framework for end-to-end autonomous driving with counterfactual-centric dataset and LLM-agents. With OmniDrive,
  1. (1)
    We introduce a counterfactual-based 3D driving Q&A design pipeline that allows for scalable, high-quality data generation.
  2. (2)
    Models pre-trained on OmniDrive showed significant improvement when tested on DriveLM Q&A benchmark and nuScenes open-loop planning, demonstrating the effectiveness and quality of our dataset.
  3. (3)
    We explore and compare two advanced frameworks Omni-L and Omni-Q, providing critical insights for designing effective LLM-Agents.
2 OmniDrive
Figure 2: The proposed counterfactual-based synthetic Q&A data generation pipeline integrates semantic key-frame selection, counterfactual-based checklist and prompt design, and human-in-the-loop quality checks to create high-quality Q&A pairs.
@@ 2 OmniDrive
We propose OmniDrive, the counterfactual-based synthetic data on nuScenes [3] with high quality Q&A pairs covering perception, reasoning and planning in 3D domain.
OmniDrive features a human-in-the-loop Q&A generation pipeline using rule-based checklist and GPT-4. As shown in Fig. 2, the data generation process can be divided into: planning-oriented key-frame selection, counterfactual-based checklist and prompt design, human-in-the-loop quality assurance, and large-scale data iteration.
2.1 Planning-oriented key-frame selection.
Autonomous driving datasets often contain significant redundancy, so we focus on selecting representative key-frames to prototype our data processing strategy. We begin by extracting CLIP [39] embeddings from the front view images of the nuScenes [3] dataset to capture diverse perceptual elements such as landmarks, traffic lights, and lane markings. Using these embeddings, we apply the K-means algorithm to cluster the data, selecting 20% of the cluster centers. This ensures that the most semantically representative data is chosen, covering various static and dynamic traffic elements. Next, we further filter the data based on the vehicle’s future trajectory. We apply K-means clustering again, this time selecting 200 cluster centers. These centers represent different vehicle dynamics, reflecting driving behaviors such as stopping, moving forward, turning left, turning right, U-turns, accelerating, decelerating, and maintaining constant speed.
This approach effectively compresses the dataset, ensuring that our algorithm design can comprehensively cover all these scenarios. By selecting keyframes, we streamline the process, allowing for more effective rule-based checklist design and prompt iteration. Once the checklist and prompts are verified to cover these scenarios, we iterate on the dataset at a larger scale. This targeted approach ensures that the most relevant data is used for further development.
2.2 Counterfactual checklist and prompt design.
Although GPT-4 has powerful ablitity in long-context processing, it cannot effectively understand 3D scenes when directly inputting images, 3D objects, lane markings, and scene definitions. It’s difficult to determine the driving status of vehicles or the relationships between traffic elements. This issue becomes more severe as the number of traffic elements in the scene increases.
To address this, we draw inspiration from counterfactual reasoning and represented the entire scene centered around simulated planned trajectories. In this section, we primarily introduce the types of prompts we input into GPT-4 and how we design checklist on counterfactual principles to enhance the quality of Q&A data generation.
Simulated trajectories. We first cluster the driving trajectories from the entire nuScenes dataset. We then classified the cluster centers into categories such as stopping, moving forward, turning left, turning right, making U-turns, accelerating, decelerating, and maintaining a constant speed. In each scene, we simulated these driving behaviors to assess their feasibility. and designed a checklist to determine whether these trajectories violated any traffic rules.
Counterfactual checklist. For fixed categories, such as object collisions, road boundary collisions, and running red lights, we use the 3D object detection, centerline and road element topology annotations from nuScenes [3] and OpenLane-v2 [45] dataset. We design a rule-based checklist to validate these scenarios.
However, relying solely on annotated perception elements cannot cover all traffic rules. Therefore, we convert the simulated driving trajectories into high-level decision-making information (this step also involves rule design, such as object and lane assignment, and determining lane-changing behavior, etc.). We then use GPT-4 to analyze the images and assess whether the driving behavior is safe and complies with traffic regulations. We found that using GPT-4 for counterfactual reasoning based on the high-level decision making still achieves good accuracy and interpretability.
Expert trajectory. We also take the log replay trajectory from nuScenes [3] as input prompt. The expert trajectories are classified into different types for high-level decision making. We also identify an object as “close”, if its minimum distance to the trajectory is smaller than 10 meters in the next 3 seconds. The close objects are then listed below the expert trajectory.
Caption. To improve the quality of Q&A data generation, we utilize counterfactual principles to structure and simplify the 3D perception annotations, avoiding the need to input lengthy and unordered scene information. To provide additional contextual information, we also prompt GPT-4 to generate captions, enhancing OCR capabilities and the recognition of open-world object categories.
When both the image and extensive scene information are fed into GPT-4 simultaneously, it tends to overlook details in the image. Therefore, we first prompt GPT-4 to produce a scene description based on multi-view input only. As shown in the top block of Tab. 1, we stitch the three frontal views and three rear views into two separate images and feed them into GPT-4. We prompt GPT-4 to include the following details: 1) mention weather, time of day, scene type, and other image contents; 2) understand the general direction of each view (e.g. the first frontal view being front-left); and 3) avoid mentioning the contents from each view independently, instead describing positions relative to the ego vehicle.
@@
The images depict a daytime setting in a controlled-access area, likely a parking lot or a service entrance of a commercial or industrial facility. On the left, there’s a grassy area with trees and a building with blue accents. Moving towards the center, we see a security checkpoint…<omitted>
Prompt type 2: Simulated decision and trajectory
@@
Prompt type 3: Expert decision and trajectory
@@
Objects need attention:
|— straight lane [(-2.6, +0.5), (+1.2, +0.7), (+5.0, +0.9), (+8.8, +1.0)]
| |— movable_object.trafficcone at (+8.2, +2.4)
| |— human.pedestrian.moving at (+4.4, -7.2)
…<omitted>
@@
Answer: The most suitable trajectory to follow would be [PT, (+0.76, +0.02), (+1.45, +0.03),…,(+3.44, +0.12)]. This trajectory ensures that you remain in your current straight lane, avoid any obstacles…<omitted>
@@
Table 1: An example to illustrate the counterfactual-based data generation pipeline. The top block shows contexts such as captions and boxes used to prompt GPT-4, and the bottom block shows the four types of responses.
2.3 Q&A generation.
In the bottom block of the Tab. 1, we describe the different types of Q&A responses obtained by using the above context information:
@@
Attention. Given the simulated and expert trajectories, run simulation to identify close objects. At the same time, we also allowed GPT-4 to use its own common sense to identify threatening traffic elements.
@@
Decision making and planning. We present the high-level decision making as well as the expert trajectory and use GPT-4 to reason why this trajectory is safe, given the previous prompt and response information as context.
@@
We design checklists on selected keyframes, followed by prompt design and Q&A generation. We manually verify the quality of the Q&A generated from these data. Once our design meets the generalization requirements, we initiate large-scale data generation. This process involves human-in-the-loop quality assurance and large-scale data iteration.
3 OmniDrive-agent
Figure 3: Overall pipeline of Omni-L and Omni-Q. The Omni-L. Omni-L follows the 2D VLM design of LLAVA, introducing 3D positional encoding and using MLP layers for vision-language alignment. Omni-Q is based on 3D BEV perception, aligning its architecture with the Q-Former design.
Designing effective Driving VLMs is a complex challenge. We consider two primary design approaches:
To explore these approaches, we propose two frameworks: Omni-L and Omni-Q. Omni-L leverages the MLP from LLaVA [25] to align multi-view image features to language embedding. Omni-Q is inspired by the BEV architecture of StreamPETR [46] and incorporates Q-Former’s [20] design to build the interaction between LLMs and traditional autonomous driving tasks.
Both Omni-L/Q use a shared visual encoder to extract multi-view image features $F_{m}\in\mathbb{R}^{N\times C\times H\times W}$. The extracted features are combined with the positional encoding $P_{m}$ and then fed into the projector. The visual features are aligned with the text in the projector and then fed into the large language model for text generation tasks. The main difference between Omni-L and Omni-Q lies in the design of the projector. One prioritizes vision-language alignment, while the other focuses on 3D perception tasks.
3.1 Omni-Q
The Transformer decoder in Q-Former [20] and the sparse query-based 3D perception models, represented by StreamPETR [46], share highly similar architecture designs. To enhance the localization abilities of the VLMs, we consider introducing the design of 3D position encoding $P_{m}$ and the supervision of the query-based perception models to the training of VLMs. As shown in Fig. 3, in QFormer, we initialize the detection queries and carrier queries and perform self-attention to exchange their information, which can be summarized by the following formula:
$\tilde{Q}=\text{Multi-head Attention}(Q,K,V).$